# Learning Graph Quantized Tokenizers for Transformers

### v1

@@ 4.2 Training
The GNN encoder is trained through gradient descent to minimize a loss function comprising three terms, where $\beta$ is the loss weight:
@@
The first term is the DGI objective, which maximizes mutual information (MI) between node representations and graph (sub-graph) representations, based on the Jensen-Shannon divergence between the joint and product of marginals as follows (Veličković et al., 2019):
$\mathcal{L}_{\text{dgi}}=\mathbb{E}\left(\sum\limits_{v\in g}\log\left(\mathcal{D}\left(h_{v},h_{g}\right)\right)+\sum\limits_{u\in\tilde{g}}\log\left(1-\mathcal{D}\left(\tilde{h}_{u},h_{g}\right)\right)\right)$ (7)
where $h_{u}$ is the representation of node $u$, $h_{g}$ is the patch (graph/sub-graph) representation that the node belongs to, $\mathcal{D}\left(.,.\right)$ is a discriminator computing the probability scores between local and global information, and $\tilde{g}$ is the corrupted version of the original graph providing negative examples. Following (Veličković et al., 2019), we define the discriminator as a bilinear layer $\mathcal{D}\left(h_{u},h_{g}\right)=\sigma\left(h_{u}^{T}\textbf{W}h_{g}\right)$, compute the global representation as a mean of node representations: $h_{g}=\frac{1}{|\mathcal{V}|}\sum\limits_{v\in g}h_{v}$, and define $\tilde{g}$ as a graph with the same structure but randomly shuffled features.
@@
where $\tilde{g}$ is the masked graph, $\tilde{h}_{v}$ is the node representation of a masked node learned by the noisy student, $h_{v}$ is the corresponding node representation learned by the teacher over the original graph, and $\gamma\geq 1$ is a scaling factor. Note that the teacher’s parameters are updated using an exponential moving average (EMA) of the noisy student’s parameters.
@@ 5 Graph Transformer
5.1 Sequence Generation
Once the tokenizer is trained, each node $v\in\mathcal{V}$ is mapped to a set of $c$ tokens: $T_{v}=\left[t^{v}_{1},\cdots,t^{v}_{c}\right]\in\mathbb{N}^{c}$, which compress information about local interactions. To enable the Transformer to capture long-range interactions, the input should consist of a sequence of tokens from nodes that are likely to have long-range dependencies. To facilitate this, we first augment the graph with semantic edges denoted as $\mathcal{E}_{s}$, which are computed as follows:
$\mathcal{E}_{s}=\left\{e_{u,v}\mid\mathop{\mathrm{arg\,topk}}\limits_{u\in\mathcal{V}}\text{sim}\left(f\left(x_{u}\right),f\left(x_{v}\right)\right)\forall v\in\mathcal{V}\right\}$ (11)
where $\text{sim}(\cdot,\cdot)$ denotes the similarity function, $x_{u}$ is the feature vector of node $u$, and $f$ is a projection function. We use cosine similarity as the similarity function and principal component analysis (PCA) as the projection function. This semantic edge augmentation effectively creates sparse edges between each node and its k-nearest neighbors in the feature space, enhancing the model’s ability to recognize and utilize significant long-range dependencies.
We then merge the semantic edges with the original graph edges and use Personalized PageRank (PPR) to generate a sequence per node. A PPR vector for a node $u$ captures the relative importance of other nodes with respect to node $u$ by exploring the graph structure through iterative random walks:
$r=\alpha\mathbf{P}r+(1-\alpha)q$ (12)
where $\mathbf{P}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}\in\mathbb{R}^{n\times n}$, $q$ is a stochastic personalized vector, $r$ is the stationary distribution of random walks, and $\alpha$ is a damping factor.
Using PPR enriches the sequence with information beyond local interactions, allowing the Transformer to access potential long-range dependencies. We construct the sequence $S_{v}$ for each node $v$ as follows:
$S_{v}=\left[T_{v}\|T_{u}\|_{u\in\mathop{\mathrm{arg\,topk}}\text{PPR}\left(v,\mathcal{E}\cup\mathcal{E}_{s}\right)}\right]$ (13)
where $S_{v}=\left[t^{v}_{1}\cdots t^{v}_{c}\mid t^{u_{1}}_{1}\cdots t^{u_{1}}_{c}\mid\cdots\mid t^{u_{k}}_{1}\cdots t^{u_{k}}_{c}\right]$ is the sequence of sorted integer tokens with length $c\times(k+1)$, based on the PPR scores for node $u$. Note that the computation of semantic edges and PPR sequences is performed only once as a pre-processing step, which reduces computational overhead during training.
@@ 5.2 Token Modulation
There are $c\times K$ possible integer tokens in total, where $c$ is the number of codebooks and $K$ is the codebook size. We randomly initialize an embedding matrix $\textbf{X}_{T}\in\mathbb{R}^{c\times K\times d_{x}}$, which is trained end-to-end with the Transformer. To further enrich the token representation, we introduce an additional token to each node by aggregating the embeddings of its assigned codebooks:
@@
We found that adding this explicit aggregated token leads to better performance compared to initializing $\textbf{X}_{T}$ with C. The input representation of the sequence for node $v$ is then defined as:
@@
This representation combines the individual token embeddings with the aggregated codebook embeddings, providing a more comprehensive and nuanced input to the Transformer.
In order to provide the Transformer with the global structural importance scores of the nodes within the sequence with respect to the target node, we introduce a gating mechanism over the input token embeddings as follows:
$S_{v}=S_{v}\odot\text{Softmax}\left(\mathop{\mathrm{topk}}\text{PPR}\left(v,\mathcal{E}\cup\mathcal{E}_{s}\right)\right)$ (16)
@@
We also introduce two trainable positional encodings to the input tokens. The first positional encoding enables the Transformer to distinguish between tokens from different nodes, while the second encoding, referred to as hierarchical encoding, allows the Transformer to recognize the hierarchy level of each token within the codebooks. We randomly initialize the positional encodings $\textbf{PE}\in\mathbb{R}^{(k+1)\times d_{x}}$ and $\textbf{HE}\in\mathbb{R}^{c\times d_{x}}$ and sum them with the encoding of their corresponding token. For example, the final encoding of the token $j$ of the node $i$ within the sequence is computed as: $x=\textbf{X}_{T}[j,t^{u_{i}}_{j}]+\textbf{PE}[i]+\textbf{HE}[j]$. Note that we did not use any structural encoding, such as Laplacian eigenvectors, as our experiments did not show any significant benefits from including them.
5.3 Transformer Encoder &\&& Classification Head
We use $l$ layers of standard Transformer encoder with flash attention (Dao et al., 2022b) to generate contextual representations per token in the sequence: $\textbf{H}^{(l)}\in\mathbb{R}^{(c+1)\times(k+1)\times d_{h}}$. We then aggregate the token representations for $j$-th node in the sequence by summing along the token dimension:
@@ 5.3 Transformer Encoder &\&& Classification Head
We feed the resulting representation into a fully-connected classifier and train the model end-to-end with cross-entropy loss. Note that during inference, only the Transformer and classifier are utilized, as the tokenizer is pretrained and the sequences are pre-computed. Furthermore, since we only require discrete tokens and codebook embeddings, our approach allows for efficient memory usage, regardless of graph size enable efficient training and inference on large-scale graphs.
@@ 6 Experiments
We comprehensively evaluate GQT on both medium-scale and large-scale node classification tasks, encompassing both homophilous and heterophilous settings across 18 datasets. Homophilous graphs are characterized by nodes with similar classes being connected to each other, whereas heterophilous graphs exhibit connections between nodes with different classes. Following the convention of most existing works on GTs, we focus on node classification tasks in our experiments. However, as discussed in Section 3, our model can be easily extended to graph classification and link prediction tasks. For each evaluation scenario, we adhere to the established experimental protocols from previous works to ensure fair comparisons. Detailed descriptions of the datasets are provided in Appendix A and detailed experimental setup and hyperparameters are provided in Appendix B.
@@ 6.1 Comparison with State-of-the-Art
Table 1: Mean node classification performance on medium-scale homophilous datasets over five runs.
@@
Measure Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$ Accuracy $\uparrow$
@@
Homophilous Node Classification. To evaluate the performance on medium-scale homophilous graphs, we use eight benchmark datasets including CoraFull (Bojchevski & Günnemann, 2017), CiteSeer, and PubMed (Yang et al., 2016), Amazon Computers, Amazon Photos, Co-author CS, and Co-author Physics (Shchur et al., 2018), as well as WikiCS (Mernyei & Cangea, 2020). We compare our results with a comprehensive set of baselines, including four traditional GNNs: GCN (Kipf & Welling, 2017), GAT (Veličković et al., 2019), APPNP (Gasteiger et al., 2018), and GPRGNN (Chien et al., 2020); four scalable GNN variants including GraphSAINT (Zeng et al., 2019), GraphSAGE (Hamilton et al., 2017b), PPRGo (Bojchevski et al., 2020), and GTAND+(Feng et al., 2022); four standard GTs including GT (Dwivedi & Bresson, 2020), Graphormer (Ying et al., 2021b), SAN (Kreuzer et al., 2021), and GraphGPS (Rampášek et al., 2022); and six state-of-the-art scalable GTs including GOAT (Kong et al., 2023a), NodeFormer (Wu et al., 2022a), DiffFormer (Wu et al., 2023a), NAGphormer (Chen et al., 2023), Exphormer (Shirzad et al., 2023), and VCR-Graphormer (Fu et al., 2024). The baseline performance is reported from existing works (Wu et al., 2023b; Luo et al., 2024a; Fu et al., 2024). As shown in Table 1, GQT outperforms the baseline GNN and GT models on 7 out of 8 benchmarks. Notably, this achievement comes with a significant reduction in memory requirement for node features during Transformer training and inference. For example, on the Physics dataset with 34,493 nodes, we only use $256\times 6$ tokens, i.e., 23-fold memory reduction.
@@
Heterophilous Node Classification. Furthermore, we evaluate GQT on six small or medium-scale heterophilous datasets: Squirrel and Chameleon (Rozemberczki et al., 2021), Questions, Roman-Empire, Amazon-Ratings, and Minesweeper (Platonov et al., 2023b). We compare the performance with seven variants of GNNs including GCN, GraphSAGE, GAT, GPRGNN, H2GCN (Zhu et al., 2020a), CPGNN (Zhu et al., 2021), and GloGNN (Li et al., 2022), and six variants of GTs, including GraphGPS, GOAT, NodeFormer, SGFormer, NAGphormer, and Exphormer. The baseline performance is reported from existing works (Wu et al., 2023b; Luo et al., 2024b; Platonov et al., 2023a; Behrouz & Hashemi, 2024). As shown in Table 2, GQT outperforms the baselines on five out of six datasets. We observe that introducing semantic edges and structural gating mechanisms specifically benefits the heterophilous setting (see Appendix C), as they enable the Transformer to capture long-range dependencies that are not easily accessible through the original graph structure.
@@
Large-scale Node Classification We also evaluate the performance of GQT on four large-scale datasets: ogbn-proteins, ogbn-arxiv, ogbn-products (Hu et al., 2020a), and pokec (Leskovec & Krevl, 2014), the last of which is a heterogeneous dataset. We compare the performance against six GNN variants: LINKX (Lim et al., 2021), SIGN (Frasca et al., 2020), GCN, GAT, GraphSAGE, and GPRGNN; and six GT variants: GraphGPS, GOAT, NodeFormer, NAGphormer, Exphormer, and SGFormer (Wu et al., 2024). We report the baseline performance from existing works (Wu et al., 2023b; Luo et al., 2024a). The results (Table 3) show that GQT outperforms the baseline models on all large-scale benchmarks. This is achieved while significantly reducing the required memory. For instance, on the ogbn-products dataset with 2,449,029 nodes and 100-dimensional node features, GQT requires only 3 codebooks of size 4096 each to represent the tokens, resulting in a remarkable 30-fold reduction in memory usage.
@@
Table 2: Mean node classification performance on heterophilous graphs over five runs.
@@ 6.2 Ablation Study
Effect of Tokenization. We examine the performance of the tokenizer by training a linear model on the representations of the learned tokens without modulation, augmentation, or Transformer (1). As shown in Table 4, within the linear evaluation protocol, the tokenizer shows strong performance, surpassing that of GTs such as GraphGPS and NAGphormer, as well as GNNs like GAT and SIGN (Table 3). This implies that the tokenizer is capable of learning effective token representations. To further investigate the importance of the tokenizer, we exclude it and train the Transformer directly on the original node features (2). As expected, this results in significant degradation in performance, highlighting the crucial role of the tokenizer. Additionally, to study the effects of vector quantization, GraphMAE2, and DGI objectives, we train the model by excluding each component (3-5). The results suggest that the SSL objectives contribute more significantly to the performance compared to vector quantization. This is because the primary purpose of vector quantization is to compress information into discrete tokens, reducing memory requirements. Between GraphMAE2 and DGI, GraphMAE2 yields the highest gain. The is due to its composition of two objectives: masked reconstruction and teacher-(noisy)student distillation. Both of these objectives have been shown to outperform InfoMax objectives on downstream tasks (Hou et al., 2022; Thakoor et al., 2022).
@@
Effect of Modulation. We also investigate the impact of codebook embeddings, positional encoding, and structural gating on the model’s performance (6-8). As shown in Table 4, introducing aggregated codebook embeddings leads to improved downstream performance due to the fact that it provides the Transformer with richer representations of each token. Positional encoding, as observed in other domains, contributes moderately to downstream performance. We also note that introducing structural gating yields moderate improvements in homophilous settings, whereas the gains are significant in heterophilous benchmarks (C). This disparity can be attributed to the ability of structural gating to provide the Transformer with importance scores computed over the global graph structure, which is particularly beneficial in heterophilous scenarios.
@@
Effect of Augmentation. We study the effect of semantic edges on downstream performance (9). The results suggest that augmenting the graph structure with semantic edges yields significant gains. This is because introducing semantic edges allows the Transformer to access semantic information that may not be captured by the original graph structure. Furthermore, when combined with random walks, this also enables the Transformer to attend to long-range dependencies. This is especially important in heterophilous benchmarks, where semantic relationships between nodes are more nuanced.
@@
(1) Linear 71.97
(2) Transformer 70.50
(3) Transformer 72.84
(4) Transformer 71.79
(5) Transformer 72.71
(6) Transformer 71.28
(7) Transformer 72.69
(8) Transformer 73.02
(9) Transformer 72.61
(10) Transformer 73.14

### v2

@@ 4.2 Training
To capture different aspects of information, we employ a multi-task learning framework that leverages three distinct families of graph self-supervised objectives: student-teacher distillation (Thakoor et al., 2022), masked autoencoding (Hou et al., 2022), and Infomax (Veličković et al., 2019). We also introduce a commitment loss (Van Den Oord et al., 2017) to enforce alignment between learned node representations and the codebook representations. Specifically, the GNN encoder is trained through gradient descent to minimize a loss function comprising of three terms, where $\beta$ is the loss weight:
@@
The first term is the Deep Graph Infomax (DGI) (Veličković et al., 2019) objective, which maximizes mutual information (MI) between node representations and graph (sub-graph) representations, based on the Jensen-Shannon divergence between the joint and product of marginals as follows:
$\mathcal{L}_{\text{dgi}}=\mathbb{E}\left(\sum\limits_{v\in g}\log\left(\mathcal{D}\left(h_{v},h_{g}\right)\right)+\sum\limits_{u\in\tilde{g}}\log\left(1-\mathcal{D}\left(h_{u},h_{g}\right)\right)\right)$ (6)
where $h_{u}$ is the representation of node $u$. $h_{g}$ is the global (sub-graph/graph) representation, computed as the mean of node representations. $\tilde{g}$ is the corrupted version of the original graph, with the same structure but randomly shuffled features, providing negative examples for contrastive learning. $\mathcal{D}\left(h_{u},h_{g}\right)=\sigma\left(h_{u}^{T}\textbf{W}h_{g}\right)$ is the discriminator that scores whether a node belongs to the graph, and is defined as a bilinear classifier.
@@
where $\tilde{g}$ is the masked graph, $\tilde{h}_{v}$ is the node representation of a masked node learned by the noisy student, $h_{v}$ is the corresponding node representation learned by the teacher over the original graph, and $\gamma\geq 1$ is a scaling factor. The teacher’s parameters are updated using an Exponential Moving Average (EMA) of the noisy student’s parameters.
@@ 5 Graph Transformer
5.1 Graph Serialization
Once the tokenizer is trained, each node $v\in\mathcal{V}$ is mapped to $c$ discrete tokens: $T_{v}=\left[t^{v}_{1},\cdots,t^{v}_{c}\right]\in\mathbb{N}^{c}$ (i.e., $T_{v}=\textbf{X}_{Q}[v]$), encoding local interactions of that node. We then need to serialize the graph in order to input it to the Transformer.
Semantic Edges. To enable the Transformer to capture long-range interactions, the input should consist of a sequence of tokens from nodes that are likely to have long-range dependencies. To facilitate this, we first augment the graph with semantic edges denoted as $\mathcal{E}_{s}$, which are computed as follows:
[…]
where $\text{sim}(\cdot,\cdot)$ denotes the similarity function, $x_{u}$ is the feature vector of node $u$, and $f$ is a projection function. We use cosine similarity as the similarity function and Principal Component Analysis (PCA) as the projection function. The semantic edge augmentation effectively creates sparse edges between each node and its k-nearest neighbors in the feature space, enhancing the model’s ability to recognize and utilize long-range dependencies.
Structural Serialization. We combine the semantic edges with the original edges and use Personalized PageRank (PPR) to generate a sequence per target node. This enriches the sequence with information beyond local interactions, allowing the Transformer to access potential long-range dependencies. We construct the sequence $S_{v}$ for each node $v$ as follows:
[…]
where $S_{v}=\left[t^{v}_{1}\cdots t^{v}_{c}\mid t^{u_{1}}_{1}\cdots t^{u_{1}}_{c}\mid\cdots\mid t^{u_{k}}_{1}\cdots t^{u_{k}}_{c}\right]$ is a sequence of length $c\times(k+1)$, comprising discrete tokens that represent the target node $v$, followed by discrete tokens of the top-$k$ relevant nodes to node $v$. These relevant nodes are determined based on PPR scores. Note that the computation of semantic edges and PPR sequences is performed only once as a pre-processing step, thereby reducing the computational overhead during training.
@@ 5.2 Token Modulation
Token Embeddings. There are $c\times K$ possible discrete tokens, where $c$ is the number of codebooks and $K$ is the codebook size. We randomly initialize an embedding matrix $\textbf{X}_{T}\in\mathbb{R}^{c\times K\times d_{x}}$, which is trained end-to-end alongside the Transformer. To further enrich the token representations, we introduce an additional token for each node that aggregating the embeddings of its assigned codebooks from the pretrained tokenizer:
@@
where $\textbf{C}[i,j]$ is the embedding corresponding to index j in the ith codebook. We found that adding this explicit aggregated token from the codebook leads to better performance compared to initializing $\textbf{X}_{T}$ directly with C. The input representation of the sequence for node $v$ is then defined as:
@@
[…]
Structural Gating. In order to provide the Transformer with the global structural importance scores of the nodes within the sequence with respect to the target node, we introduce a gating mechanism over the input token embeddings as follows:
[…]
@@
Positional Encoding. We also introduce two trainable positional encodings to the input tokens. The first positional encoding enables the Transformer to distinguish between tokens from different nodes, while the second encoding, referred to as hierarchical encoding, allows the Transformer to recognize the hierarchy level of each token within the codebooks. We randomly initialize the positional encodings $\textbf{PE}\in\mathbb{R}^{(k+1)\times d_{x}}$ and $\textbf{HE}\in\mathbb{R}^{c\times d_{x}}$ and sum them with the encoding of their corresponding token. For example, the final encoding of the token $j$ of the node $i$ within the sequence is computed as: $x=\textbf{X}_{T}[j,t^{u_{i}}_{j}]+\textbf{PE}[i]+\textbf{HE}[j]$. Note that we did not use any structural encoding, such as Laplacian eigenvectors, as we did not observe any significant gains from them.
5.3 Transformer Encoder
We use $l$ layers of standard Transformer encoder with flash attention (Dao et al., 2022) to generate contextual representations per token in the sequence: $\textbf{H}^{(l)}\in\mathbb{R}^{(c+1)\times(k+1)\times d_{h}}$. We then aggregate the token representations for $j$-th node in the sequence by summing along the token dimension:
@@ 5.3 Transformer Encoder
We feed the resulting representation into a fully-connected classifier and train the model end-to-end using cross-entropy loss. Note that during inference, only the Transformer and classifier are utilized, as the tokenizer is pretrained and the sequences are pre-computed. Furthermore, since we only require discrete tokens and codebook embeddings, our approach enables efficient memory usage, regardless of graph size, allowing for efficient training and inference on large-scale graphs.
@@ 6 Experiments
We evaluate GQT on both medium- and large-scale graph learning tasks, encompassing 22 homophilic, heterophilic, and long-range benchmarks. We follow the established experimental protocols from previous works to ensure fair comparisons. Details of the datasets, experimental setup, and hyperparameters are provided in Appendices 10 and 11, respectively.
@@ 6.1 Comparison with State-of-the-Art
Long-Range Benchmarks. We use four datasets from the Long-Range Graph Benchmark (LRGB) (Dwivedi et al., 2022b), including the Peptides-Func dataset for graph classification with Average Precision (AP) metric, the Peptides-Struct dataset for graph regression with Mean Absolute Error (MAE) metric, the COCO-SP dataset for inductive node classification with macro F1 metric, and the PCQM-Contact for link prediction with Mean Reciprocal Rank (MRR) metric. We compare our results to baselines reported in (Wang et al., 2024). The results shown in Table 1 suggest that GQT is able to capture long-range dependencies and performs well on various graph prediction tasks.
Table 1: Mean performance on inductive long-range benchmarks over five runs.
Task Graph Classification Graph Regression Node Classification Link Prediction
Dataset Peptides-Func Peptides-Struct COCO-SP PCQM-Contact
#Graphs 15,535 15,535 123,286 529,434
Avg. #Nodes 150.94 150.94 476.88 30.14
Avg. #Edges 307.30 307.30 2,693.67 61.09
Metric AP ↑ MAE ↓ F1 ↑ MRR ↑
GCN 0.5930$\pm$0.0023 0.3496$\pm$0.0013 0.0841$\pm$0.0010 0.3234$\pm$0.0006
Exphormer 0.6258$\pm$0.0092 0.2512$\pm$0.0025 0.3430$\pm$0.0108 0.3587$\pm$0.0025
GPS 0.6535$\pm$0.0041 0.2500$\pm$0.0005 0.3412$\pm$0.0044 0.3337$\pm$0.0006
Graph-Mamba 0.6739$\pm$0.0087 0.2478$\pm$0.0016 0.3960$\pm$0.0175 0.3395$\pm$0.0013
GQT (Ours) 0.6903$\pm$0.0085 0.2452$\pm$0.0018 0.4007$\pm$0.0125 0.3427$\pm$0.0012
Table 2: Mean node classification accuracy on medium-scale homophilic datasets over five runs.
@@
Homophilic Node Classification. We use eight medium-scale homophilic datasets including: CoraFull (Bojchevski and Günnemann, 2017), CiteSeer, PubMed (Yang et al., 2016), Amazon Computers, Amazon Photos, Co-author CS, Co-author Physics (Shchur et al., 2018), and WikiCS (Mernyei and Cangea, 2020). We compare our results with eight GNNs including: GCN (Kipf and Welling, 2017), GAT, APPNP (Gasteiger et al., 2018), GPRGNN (Chien et al., 2020), GraphSAINT (Zeng et al., 2020), GraphSAGE (Hamilton et al., 2017), PPRGo (Bojchevski et al., 2020), and GTAND+ (Feng et al., 2022). We also compare against ten GTs including GT (Dwivedi and Bresson, 2020), Graphormer (Ying et al., 2021), SAN (Kreuzer et al., 2021), GraphGPS (Rampášek et al., 2022), GOAT (Kong et al., 2023), NodeFormer (Wu et al., 2022), DiffFormer (Wu et al., 2023a), NAGphormer (Chen et al., 2023), Exphormer (Shirzad et al., 2023), and VCR-Graphormer (Fu et al., 2024). The baseline performance is reported from (Wu et al., 2023b; Luo et al., 2024a). GQT outperforms the baseline GNN and GT models on 7 out of 8 benchmarks (Table 2). Notably, this achievement comes with a significant memory reduction. For example, on the Physics dataset with 34,493 nodes, we only use $256\times 6$ tokens, i.e., a 23-fold memory reduction.
@@
Heterophilic Node Classification. We also evaluate GQT on six medium-scale heterophilic datasets: Squirrel, Chameleon (Rozemberczki et al., 2021), Questions, Roman-Empire, Amazon-Ratings, and Minesweeper (Platonov et al., 2023). We compare the performance with seven GNNs: GCN, GraphSAGE, GAT, GPRGNN, H2GCN (Zhu et al., 2020a), CPGNN (Zhu et al., 2021), and GloGNN (Li et al., 2022), and six GTs: GraphGPS, GOAT, NodeFormer, SGFormer, NAGphormer, and Exphormer. The baseline performance is reported from (Wu et al., 2023b; Luo et al., 2024b; Platonov et al., 2023; Behrouz and Hashemi, 2024). As shown in Table 3, GQT outperforms the baselines on five out of six datasets. We observe that introducing semantic edges and structural gating specifically benefits the heterophilic setting (Appendix 12.1), as they enable the Transformer to capture long-range dependencies that are not easily accessible through the original graph structure.
@@
Large-scale Node Classification We also use four large-scale datasets: ogbn-proteins, ogbn-arxiv, ogbn-products (Hu et al., 2020a), and pokec (heterogeneous) (Leskovec and Krevl, 2014). We compare the performance against six GNN: LINKX (Lim et al., 2021), SIGN (Frasca et al., 2020), GCN, GAT, GraphSAGE, and GPRGNN; and six GTs: GraphGPS, GOAT, NodeFormer, NAGphormer, Exphormer, and SGFormer (Wu et al., 2023b). We report the baseline performance from (Wu et al., 2023b; Luo et al., 2024a). The results (Table 4) show that GQT outperforms the baseline models on all large-scale benchmarks. This achievement comes with a significant reduction in required memory. For instance, on the ogbn-products dataset with 2,449,029 nodes and 100-dimensional node features, GQT requires only 3 codebooks of size 4096, resulting in a 30-fold memory reduction.
@@
Table 3: Mean node classification performance on heterophilic graphs over five runs.
@@ 6.2 Ablation Study
Effect of Tokenization. We examine the performance of the tokenizer by training a linear model on the representations of the learned tokens without modulation, augmentation, or Transformer (1). As shown in Table 6, within the linear evaluation protocol, the tokenizer shows strong performance, surpassing that of GTs such as GraphGPS and NAGphormer, as well as GNNs like GAT and SIGN (Table 4). This implies that the tokenizer is capable of learning effective token representations. To further investigate the importance of the tokenizer, we exclude it and train the Transformer directly on the original node features (2). As expected, this results in significant degradation in performance, highlighting the crucial role of the tokenizer. Additionally, to study the effects of vector quantization, GraphMAE2, and DGI objectives, we train the model by excluding each component (3-5). The results suggest that the SSL objectives contribute more significantly to the performance compared to vector quantization. This is because the primary purpose of vector quantization is to compress information into discrete tokens, reducing memory requirements. Between GraphMAE2 and DGI, the former yields the highest gain. This is due to its composition of two objectives: masked reconstruction and teacher-(noisy)student distillation. Both of these objectives have been shown to outperform InfoMax objectives on downstream tasks (Hou et al., 2022; Thakoor et al., 2022).
@@
Effect of Modulation. We also investigate the impact of codebook embeddings, positional encoding, and structural gating on the model’s performance (6-8). As shown in Table 6, introducing aggregated codebook embeddings leads to improved downstream performance because it provides the Transformer with richer representations of each token. Positional encoding, as observed in other domains, contributes moderately to downstream performance. We also note that introducing structural gating yields moderate improvements in homophilic settings, whereas the gains are significant in heterophilic benchmarks (12.1). This disparity can be attributed to the ability of structural gating to provide the Transformer with importance scores computed over the global graph structure, which is particularly beneficial in heterophilic scenarios.
@@
Effect of Augmentation. We study the effect of semantic edges on downstream performance (9). The results suggest that augmenting the graph structure with semantic edges yields significant gains. This is because introducing semantic edges allows the Transformer to access semantic information that may not be captured by the original graph structure. Furthermore, when combined with random walks, this also enables the Transformer to attend to long-range dependencies which is particularly important in heterophilic benchmarks, where semantic relationships between nodes are more nuanced.
Table 5: Accuracy drop under adversarial attack.
Attack GR-BCD PR-BCD
PubMed ogbn-arxiv PubMed ogbn-arxiv
RQ-VAE 20.40% 14.80% 23.30% 17.20%
GQT (ours) 15.80% 10.40% 18.10% 11.30%
Robustness Analysis. To measure robustness, we use Greedy Randomized Block Coordinate Descent (GRBCD) and Projected Randomized Block Coordinate Descent (PR-BCD) (Geisler et al., 2021) adversarial attacks to measure the accuracy degradation. We compare the GQT with an RQ-VAE (Lee et al., 2022). The results (Table 5) show that our tokenizer is more robust to attacks. This is because GQT is trained with multi-task self-supervised objectives while RQ-VAE is trained with reconstruction objective. More details are provided in Appendix 12.
@@
[…]